20201129 D/E

08:07 Call from Andrew Ca that swing arms not working. Will do AO

08:30 After stop_all/start_all and power on, the OCAM was not live
  • followed Fabio's trouble shooting guide with power off and process restarts but stay not live
  • ocam_error.jpg:
    ocam_error.jpg

08:38 Power off, stop_all/start_all, power on

08:53 OCAM OK Now. Cooling

08:50 SX AdSec processes disconnect from msgdd and system offline.
  • This is the known bug of cycling the WUnit kills AdSec processes

09:00 Rerotator was not moving and center pupil was driving Tip/Tilt the wrong direction
  • Canceled preset
  • Stop/start WFS Arbitrator
  • Send preset again
  • All OK

09:07 Loop Closed
  • GOPT On
  • NCPA Off
  • LUCI 3-22
  • see log
    • /home/AOeng/doug/get_telemetry/OSCO/20201130_090716.log
    • /home/AOeng/doug/get_telemetry/OSCO/20201130_091028.log

09:14 NCPA On
  • Collect another set of OSCO
  • LUCI 24-42
  • see log
    • /home/AOeng/doug/get_telemetry/OSCO/20201130_091625.log
    • /home/AOeng/doug/get_telemetry/OSCO/20201130_091812.log

09:21 Off to SL photometric standard

10:23 ESM preset
  • Set to AUTO mode to see how new ESM steps work
  • seeing 0.85"
  • presetAO failed
  • baysidex timeout
WfsArbitrator_2522 |ERR|       421|2020-11-30 10:27:39.444781|             MAIN > Timeout error on baysidex (-5001) TIMEOUT_ERROR
WfsArbitrator_2522 |DEB|       422|2020-11-30 10:27:39.445136|             MAIN > arbFunc updateModulationInternal (, 'FLAOWFS', 'LUCIFER', 2000)
WfsArbitrator_2522 |INF|       423|2020-11-30 10:27:39.445461|             MAIN > Tip-tilt frequency is not synced, skipping update

CopleyCtrl_155412  |ERR|       215|2020-11-30 10:29:11.558589|             MAIN > Error: movement out of range
CopleyCtrl_155412  |ERR|       216|2020-11-30 10:29:11.558665|             MAIN > Request: -108.666832   Limits: 0.000000 -108.000000
CopleyCtrl_155412  |ERR|       217|2020-11-30 10:29:11.558768|             MAIN > Variable handler error: [AOException] Error in MoveTo(): movement out of range (code -5002) Value is out of range
  • Ref star out of patrol field

10:43 Preset
  • Center Star
  • Close loop
  • Done, but gains are LNgains.fit
  • When loop opens, Re-close loop is not available
  • 16.9 mag

WfsArbitrator_2522 |INF|       643|2020-11-30 10:43:15.258070|             MAIN > selectHODark(): selected dark file 20201130_104313.fits
WfsArbitrator_2522 |INF|       644|2020-11-30 10:43:18.590501|             MAIN > calibHODarkInternal(): done
WfsArbitrator_2522 |DEB|       645|2020-11-30 10:43:29.085939|             MAIN > arbFunc doApplySettings done.
WfsArbitrator_2522 |DEB|       646|2020-11-30 10:43:31.802803|             MAIN > arbFunc doCloseLoop (, 'FLAOWFS', 'LUCIFER')
WfsArbitrator_2522 |DEB|       647|2020-11-30 10:43:31.804782|             MAIN > arbFunc doCloseLoop done.
WfsArbitrator_2522 |DEB|       648|2020-11-30 10:43:55.625115|             MAIN > arbFunc doPauseLoop (, 'FLAOWFS', 'LUCIFER')

11:00 Tried gains of 9.7, 12.5. Skipframe and loop opens
  • reclose loop from the AOS Command GUI
  • set gains to 2.0 1.0, 0.0

11:15 hit co-pointing limit so offset out of FOV failed
  • had to ResumeAO from AOS Command GUI before start LUCI script

12:07 AdSec went into failure during ESM closed loop. Not clear why
  • Had to fsm_power_off() then power on at AdSec Control GUI

12:30 For all of this field, LUCI 2 is better (3.5 pix FWHM) than LUCI1 (4.3, 5.2 pix FWHM) in ESM mode.
  • not sure why
  • TN 123043, 123321
  • Gains at 2.5, 1.0. Can not go much higher becasue of skipframes.

12:41 Preset to 13.5 mag ESM star
  • gains 2.0, 1.0
  • LUCI1 (3.3, 3.1 FWHM) LUCI2 (4.1 FWHM)
  • gains 5.0, 1.8
  • LUCI1 (3.2, 3.9) (2.9. 3.4)
  • TN 125102, 125406

-- DougMiller - 30 Nov 2020
Topic revision: r2 - 01 Dec 2020, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback